A random variable X has Poisson distribution with mean equal to 0.4. What is the
probability that the random variable is greater than zero?
1
Expert's answer
2020-06-01T19:01:52-0400

The aim is to compute "P(X>0)". We remind that for the Poisson distribution with the mean equal to "0.4" we have

"P(X=k)= \frac{(0.4)^{k}}{k!}e^{-0.4}."

Thus, using the Taylor series for the exponent, we have

"P(X>0)=\sum_{k=1}^{+\infty} \frac{(0.4)^{k}}{k!}e^{-0.4}=e^{-0.4}(e^{0.4}-1)=0.3297"